ASDC has a contract to produce eight satellites to support a worldwide telephone system that allows individuals to use a single, portable telephone in any location on earth to call in and out. ASDC has estimated that the R&D costs will be USD $1,000,000. Material costs are expected to be $500,000. They have estimated the design and production of the first satellite will require 100,000 labor hours and a 90 percent improvement curve is expected. Skilled labor cost is $30 per hour. Desired profit for all projects is 25 percent of total costs. Answer the following questions. You may use excel

a) How many labor hours should the eighth satellite require?

b) How many labor hours for the whole project of eight satellites?

c) What price would you ask for the project? Why?

d) Midway through the project your design and production people realize that an 85 percent improvement curve is more appropriate. What impact does this have on the project?
Answer #1

(a) Labout hours needed by nth unit

Tn = T1 * ( n)b where b = log ( learning rate) / log2

T8 = 100000 x (8)b where b =-0.152

T8 = 100000 x (8)-0.152 = 72900

(b) T1 to T8

100000 [ 1+ 0.9+(3)-0.152 + (4)-0.152 + (5)-0.152+ (6)-0.152 + (7)-0.152 + (8)-0.152]

100000x [ 1.9+ 0.8462+0.81+0.7829+0.7615+0.7439+0.729]

= 6.57355 x100000 = 657355

(c) Total cost =1.25 x  [production cost + material cost + research and develpment cost]

= 1.25 x [ 657355 x30 + 500000+1000000]

= 26525821

(d) If an 85% learning curve is expected, it means that the number fo hours expected to complete the units will be lesser than that in 90% LR. Therefore the cost estimation need to be revised and will be lesser than value with 90%LR